a mob of pigs walk towards you at 7.00m/s. When they are 42.0m from you, you lift a frying pan and they begin to accelerate towards you with a constant rate of 2.00m/s2. How fast are they moving when they get you?
Answer #1

Initially velocity u= 7 m/s Distance s=42m. az 2 mi82 Newhans third can v²_u²-zas. 2 (7)?= 2x2 x 42 uz (7)² + 4x42 w=5 (72+